the perimeter of a rectangle is 100 yards. What are the dimensions of the rectangle if the length is 40 yards more then the width?
The lendth is ? yards and the width is ? yards

Answer by rfer(16322)

P=2w+2l
2x+2(x+40)=100
2x+2x+80=100
4x=20
x=5
x+40=45